OpenStudy (anonymous):

-3 4/7 in simplest form. will fan and medal

OpenStudy (anonymous):

If that's a mixed number, then (-3/1)*(7/7) + (-4/7) = (-21/7) + (-4/7) = -25/7

OpenStudy (anonymous):

awesome! thanks

OpenStudy (anonymous):

:)